One litre of co2 is passes through red hot coke.the volume becomes 1.4 litres at same temperature and pressure. The composition of products are

Grade:11

1 Answers

Ashutosh Sharma
8 years ago

The equation can be represented as,
CO2 + C → 2CO
Now let x be litre of CO2 which combines with coke to form 2x L of CO.
Volume of CO2 left =1-x
Total volume =1.4 L
1-x+ 2x =1.4
x=0.4
Volume of CO=2×0.4=0.8 L
Volume of CO2=1- 0.4 lt=0.6 L